I'm having issues with this problem.
-7(t+2)=3(4t+1)
I'm geting -17/19?

Answer by longjonsilver(2297)

-7(t+2)=3(4t+1)
-7t-14 = 12t+3
-7t-12t = 3+14
-19t = 17
t = -17/19....have confidence, you are correct :-)

The obvious thing to do, is to check your answer in the original question:

-7(t+2) --> -7 and 14/19 (when we put t=-17/19 in it)
3(4t+1) --> -7 and 14/19 (when we put t=-17/19 in it)

So, this answer is correct.
jon